# SEA: Low-Resource Safety Alignment for Multimodal Large Language Models via Synthetic Embeddings

### v1

@@ B.1 More Implementation Details of SEA
Figure 5: Prompt for harmful phrase extraction
@@
Details for Content Control and Style Control Samples. Both samples consist of an instruction, a response prefixes, and a response. For the content control sample, the response is the sentence completed from harmful phrases. For the style control sample, the response is sampled from a pre-prepared style set. Each model’s instruction, response prefixes, and style set are different and are designed based on the output patterns of the model. Specifically, we used 50 test samples to observe each model’s output and summarize their habits. Figure 7 displays the model’s instructions, response prefixes, and style sets.
@@ B.2 More Details on Experimental Setup
Training Data Construction. We sampled 2k harmful samples and 1k harmless samples from the SafeRLHF dataset. Each sample in SafeRLHF includes an instruction, a chosen response, and a rejected response. The harmful samples were randomly selected from the rejected responses with a severity level of 3. In contrast, the harmless samples were randomly selected from those where both the chosen and rejected responses were equal to 0. Since we found that a significant amount of harmful content still existed in the chosen responses, we used Llama2-7b-chat to regenerate the chosen responses. In any SFT training, only the instruction and chosen response from each sample are used for training.
Evaluation Setting For MM-SafetyBench, we selected only the same eight scenarios for evaluation as in VA-SafetyBench. In Tables 3, 4, and 5, the average ASR for each task in MM-SafetyBench and VA-SafetyBench is presented across these eight scenarios. For all general capability evaluation benchmarks, we follow the implementation in Ji et al. (2024b).
@@
Figure 7: Instruction, response prefixes, and style sets designed for different models.
Training Setting for Safety Alignment. All methods for safety alignment training are implemented on a server equipped with four A800 GPUs. For VLGuard, we follow the original paper’s setup and retrain on the Hugging Face version of Llava-v1.5-7b-hf. For other baselines, we summarize the training hyperparameter settings in Table 10.
Table 10: Hyperparameter settings for safe alignment training. (SFT) indicates that SEA SFT and Textual SEA share the same set of hyperparameters.
@@
LLaVA-v1.5-7b-hf (SFT) full-parameter 2e-5 2
LLaVA-v1.5-7b-hf (DPO) full-parameter 2e-6 3
@@
B.3 Case study
To further validate the effectiveness of SEA, Figures 8, 9, and 10 respectively present cases evaluated on the safety benchmarks for image, video, and audio MLLMs. All cases come from the most challenging tasks in their respective modality benchmarks, namely SD+TYPO, DM+TYPO, and KPT+noise.
Figure 8: A case on the security evaluation of the image-based MLLMs. The evaluation sample is sourced from the SD+TYPO task of MM-SafetyBench, with the scenario “Sex”.
Figure 9: A case on the security evaluation of the video-based MLLMs. The evaluation sample is sourced from the DM+TYPO task of VA-SafetyBench, with the scenario “physical harm”.

### v3

@@ B.1 More Implementation Details of SEA
Figure 5: Prompt for harmful phrase extraction.
@@
Details for Content Control and Style Control Samples. Both samples consist of an instruction, a response prefixes, and a guiding text. For the content control sample, the guiding text is either a harmful phrase (for product) or a sentence completed from harmful phrase (for activity). For the style control sample, the guiding text is a style description randomly sampled from a pre-prepared style set. Each model’s instruction, response prefixes, and style set are different and are designed based on the output patterns of the model. Specifically, we used 50 test samples to observe each model’s output and summarize their habits. Figure 7 displays the instructions, response prefixes, and style sets used by the three MLLMs. For easier understanding, we also prepared an example for each MLLM, which includes all intermediate outputs generated during the preparation of the embedding optimization dataset.
@@ B.2 More Details on Experimental Setup
Training Data Construction. We sampled 2k harmful samples and 1k harmless samples from the SafeRLHF dataset. Each sample in SafeRLHF includes an instruction, a chosen response, and a rejected response. The harmful samples were randomly selected from samples in which the rejected responses with a severity level of 3. In contrast, the harmless samples were randomly selected from samples in which both the chosen and rejected responses had a sensitivity level of 0. Since we found that a significant amount of harmful content still existed in the chosen responses, we used Llama2-7b-chat to regenerate the chosen responses for harmful samples. In any SFT training, only the instruction and chosen response from each sample are used for training.
Evaluation Setting. For MM-SafetyBench, we selected only the same eight scenarios for evaluation as in VA-SafetyBench. In Tables 3, 4, and 5, the results presented by VA-SafetyBench and MM-SafetyBench are the average ASR for the eight scenarios. For all general capability evaluation benchmarks, we follow the implementation in Ji et al. (2024b).
@@
Figure 7: Guidelines for constructing embedding optimization datasets. Each MLLM includes unique instructions, response prefixes, and a style set. Examples containing all intermediate outputs is provided to aid in understanding the construction process in Section 3.2.
Training Setting for Safety Alignment. All methods for safety alignment training are implemented on a server equipped with four A800 GPUs. For VLGuard, we follow the original paper’s setup and retrain on the Huggingface version of LLaVA-1.5-7b-hf. For other baselines, we summarize the training hyperparameter settings in Table 10.
Table 10: Hyperparameter settings for safe alignment training. (SFT) and (DPO) indicate the training strategy.
@@
LLaVA-1.5-7b-hf (SFT) full-parameter 2e-5 2
LLaVA-1.5-7b-hf (DPO) full-parameter 2e-6 3
@@
B.3 The Impact of Style Control Samples
To verify whether style control samples help enhance the diversity of embeddings, we removed the style control samples and performed embedding optimization using only the content control samples. We then calculated the average cosine distance between the embeddings with and without the style control samples.
Table 11 shows the results. The consistent results across three different modal MLLMs indicate that adding style control samples enhances the average differences between embeddings, resulting in greater diversity. In addition, Figure 8 shows the t-SNE visualization of the SEA embeddings. Points of different styles in LLaVA-1.5-7b-hf and Qwen2-VL-7b form clusters in the embedding space, while in Qwen2-Audio-7b, points of different styles are distributed along a semicircular arc at different locations. It’s worth mentioning that when style optimization is not specified, we observe that the SEA embeddings obtained by LLaVA-1.5-7b-hf consistently correspond to the “black and white photograph” style. This may be related to the fact that we always use a fixed embedding (such as the embedding of a white image) as a starting point for optimization. Therefore, specifying the embedding style enables a broader distribution in the feature space, enhancing the diversity of the embeddings.
Table 11: Comparison of average cosine distances of SEA embeddings with and without style control.
Models SEA SEA without style control
LLaVA-1.5-7b-hf 0.03109 0.03023
Qwen2-VL-7b 0.06986 0.06663
Qwen2-Audio-7b 0.00796 0.00745
Figure 8: The t-SNE visualization of the SEA embeddings
B.4 Case Study
Embedding Quality. To explore what the optimized embeddings capture, we first ignored the modality encoder of MLLMs, and then input both synthetic embedding and textual instruction into the unaligned MLLM simultaneously. Each MLLM utilized three different instructions combined with the same embedding for single-round inference. The experimental results shown in Figure 9 indicate that even when the content guiding text consists of only brief descriptions, the optimized embeddings can capture rich and realistic information, such as that bombs should have fuses or the specific methods used by drug traffickers for money laundering. Furthermore, when faced with different queries, MLLM consistently interprets the embeddings as products or activities within the content guiding texts, as if these embeddings were encoded from real multimodal data containing this information.
It must be acknowledged that during the testing of some SEA embeddings, we observed hallucinations. This may be due to the model having not encountered similar data during pre-training stage, indicating a lack of relevant additional modal knowledge. In Section 5.3, we analyzed that retaining such low-quality embeddings does not compromise safety but may harm the model’s general performance. Implementing an output checking mechanism in such tests could potentially filter out low-quality embeddings, and we encourage future work to explore this.
Alignment Performance. To further explore how multimodal alignment data incorporating SEA embeddings enhances safety alignment performance, Figures 10, 11, and 12 respectively present cases evaluated on safety benchmarks for image, video and audio-based MLLMs. All cases come from the most challenging tasks in their respective modality benchmarks, namely SD+TYPO, DM+TYPO, and KPT+noise. When faced with harmful queries in three cases, both SEA SFT and SEA DPO refused to respond and provided their reasons.
Figure 9: Three examples of using SEA embeddings for dialogue with MLLMs. Each SEA embedding tested three instructions. Each instruction was fed into the MLLM along with the SEA embedding for single-round inference.
Figure 10: A case on the security evaluation of the image-based MLLMs. The evaluation sample is sourced from the SD+TYPO task of MM-SafetyBench, with the scenario “Sex.”
Figure 11: A case on the security evaluation of the video-based MLLMs. The evaluation sample is sourced from the DM+TYPO task of VA-SafetyBench, with the scenario “physical harm.”